Experiment 13: How Much Stomach Acid Cam Antacids Neutralize? PREPARATION WORKSHEET Name: Date TA: Section nswer the following questions using information obtained from lecture, the textbook, the lab manual, or other reliable sources. The responses will be collected at the beginning of Pre-Lab ation and checked at the start of your lab section for credit. Failure to complete all answers may result in exclusion from participating in the lab. 1. For each of the following food types, list where digestion is initiated and what enzymes are used to hydrolyze them. a. carbohydrates: b. proteins: c. lipids: 2. How does the pHl of the stomach change when eating versus not eating? 3. In two or three complete sentences, describe what heartburn is and how antacids work to chemically counteract heartburn.

media%2F253%2F25391b01-f4c9-4fee-99f9-91
Answer #1

CARBOHYDRATE -Carbohydrate digestion begins in the mouth. The salivary glands in the mouth secrete saliva, which helps to moisten the food. The food is then chewed while the salivary glands also release the enzyme salivary amylase, which begins the process of breaking down the polysaccharides in the carbohydrate food.

The two digestible carbohydrates are starches and sugars, and both of these carbohydrates are digested, or broken down into their most elementary form, along the gastrointestinal tract. Amylase, an enzyme which breaks apart starches, is found in the mouth and in the small intestine.

PROTINE -Pepsin is the active protein-digesting enzyme of the stomach. Pepsin acts on protein molecules by breaking the peptide bonds that hold the molecules together.Digestion of protein is completed in the small intestine by the pancreatic enzymes trypsin, chymotrypsin, and carboxypeptidase.

Lipids -Most lipids that you consume in your diet are fats. Some digestion occurs in your mouth and the stomach, but most takes place in the small intestine. Bile is produced by your liver, stored and released in your gall bladder and emulsifies fat globules into smaller droplets. This greatly increases the surface area that allows lipase, a fat-digesting pancreatic enzyme, to aid in digestion. After digestion, these broken-down fat particles called fatty acids combine with cholesterol and bile to move into your cells' mucosa where they are reconverted into large molecules, most passing into vessels -- called lymphatics -- near the intestine. These vessels transport fat to the veins of your chest, and the blood carries fat to be stored in adipose tissue throughout your body.

5) No. of moles of NaOH = volume * concentration

= 13.34 ml * .09564 mol / L

= 1.28milimol

Concentration of HCl can be calculated by M1V1 (NaOH) = M2V2(HCl)

.09564 * 13.34 = M2 * 15.13

M2 = 1.275 / 15.13

= .0843M

Now No. of MOles Of HCl = .0843* 15.13

= 1.28 milimoles (same as moles of NaOH)